OpenStudy (anonymous):

74. Tom has the following medical bills. If he has 15% insurance co-payment, how much does he pay? Hospital: $52,578 Doctor: $1,380 Lab work: $3,243

OpenStudy (anonymous):

Hospital: $52,578*(0.15) = $7886.7 Doctor: $1,380*(0.15) = $207 Lab work: $3,243*(0.15) = $486.45

OpenStudy (anonymous):

Thank you